OpenStudy (anonymous):

Which two square roots are used to estimate √42

OpenStudy (anonymous):

\[\sqrt{7}*\sqrt{6}\]

OpenStudy (anonymous):

Thats not an answer it's either 25-36 36-49 49-64 25-64

OpenStudy (anonymous):

Someone plzz help

zepdrix (zepdrix):

Is that subtraction between the options? Hmm that's confusing..

OpenStudy (anonymous):

No i meant and

zepdrix (zepdrix):

Oh :p

OpenStudy (anonymous):

Which one is it ughh I'm solo confused......... lol

zepdrix (zepdrix):

I kinda like what peach was doing, writing it as 6*7 under the root. So think about this:\[\large\rm \sqrt{6\cdot6}\lt \sqrt{6\cdot7}\lt\sqrt{7\cdot7}\]I put the smaller one on the left, the larger one on the right. So then,\[\large\rm \sqrt{36}\lt \sqrt{42}\lt\sqrt{49}\]Got it? :)

OpenStudy (anonymous):

OMG yes thank u soo much